### v3

# On the Non-Identifiability of Steering Vectors in Large Language Models

Abstract
Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations. This interpretation assumes that steering directions are identifiable and uniquely recoverable from input–output behavior. We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits. Critically, we show that non-identifiability is a robust geometric property that persists across diverse prompt distributions. These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
On the Non-Identifiability of Steering Vectors in Large Language Models
Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com        Ashish Mahendran Kurapath1 1Manipal Institute of Technology Bengaluru
1 Introduction
Persona vector steering (Chen et al., 2025) has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations. This line of work is closely connected to broader efforts in representation engineering (Zou et al., 2023) and activation editing (Rimsky et al., 2024; Turner et al., 2023), where linear directions in activation space are used to modulate model behavior along semantic axes (Elhage et al., 2022; Turner et al., 2024).
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful and uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
This raises a fundamental question: when does representational alignment afford reliable behavioral control? Understanding identifiability is critical for several reasons. First, it determines when persona vectors provide principled causal interventions versus merely heuristic control that exploits one of many behaviorally equivalent directions. Second, it clarifies when claims that a specific vector “represents” a semantic concept are scientifically well-grounded rather than artifacts of measurement and projection (Elazar et al., 2021; Marks and Tegmark, 2023). Third, it reveals when steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting. Finally, identifiability theory clarifies which experimental protocols provide meaningful evidence for causal interpretation of interventions on internal activations (Schölkopf et al., 2021; Ahuja et al., 2022).
Figure 1: Steering vector non-identifiability: The original vector $v$ and a perturbed vector $v+v_{\perp}$ with $v_{\perp}\in\mathrm{ker}(J)$ produce identical outputs since null-space components do not affect JvJv, the first-order change in output logits induced by steering.
We make three contributions. First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity. Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits. Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than just an artifact of limited prompt diversity. These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
2 Related Work
Our work formalizes persona steering [^1] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025). as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
Causal and latent variable identifiability
Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977). Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
Probing and representation learning
Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information. The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
Activation editing in LLMs
Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined. Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
Robustness and invariant representations
Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations. Methods for learning invariant representations (Arjovsky et al., 2019; Krueger et al., 2021) aim to extract features stable across domains. Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
Mechanistic interpretability
Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints. Our theoretical analysis clarifies when such directions are well-defined.
3 Problem Setup
3.1 Formal Model
Consider a pre-trained transformer language model $f_{\theta}$ with $L$ layers. For a given input prompt $x$ (tokenized as $x_{1},\ldots,x_{T}$), let $h_{\ell}(x)\in\mathbb{R}^{d}$ denote the hidden representation at layer $\ell$ and position $T$ (typically the final token position for autoregressive generation).
Latent persona variable
We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or "persona" (e.g., formality, political stance, truthfulness).
Steering intervention
A steering vector $v\in\mathbb{R}^{d}$ is applied as:
$\tilde{h}\ell(x)=h\ell(x)+\alpha v$ (1)
where $\alpha\in\mathbb{R}$ is the steering strength. The modified representation $\tilde{h}_{\ell}$ is fed forward through subsequent layers to produce output logits $o(x,v,\alpha)$ over the vocabulary.
Generative model
We posit that the true data-generating process involves:
$z\sim p(z),\quad h_{\ell}=g_{\ell}(x,z)+\epsilon$ (2)
where $g_{\ell}:\mathcal{X}\times\mathcal{Z}\to\mathbb{R}^{d}$ encodes how persona $z$ modulates the representation for prompt $x$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is measurement noise. The goal of steering is to approximate the effect of varying $z$ by adding $v$.
3.2 Observational Regimes
We consider two data access regimes that determine what alignment interventions can afford:
  • Regime 1: Black-box input–output The researcher observes only $(x,y)$ pairs, where $y$ is generated text. There is no access to internal representations. This is the weakest regime and corresponds to behavioral evaluation.
  • Regime 2: White-box single-layer access The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$. This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values (e.g., polite versus rude instructions):
$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})]$ (3)
Our analysis focuses primarily on Regime 2, as it represents the dominant paradigm in current steering research.
3.3 Linear Approximation and Nonlinear Case
Local linearization
Near a reference distribution, we can approximate the effect of steering on output logits as:
$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (4)
where $J_{\ell}(x)=\left.\frac{\partial o}{\partial h_{\ell}}\right|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ denotes the Jacobian of the output and $V$ is the vocabulary size.
Nonlinear case
In general, the mapping $h_{\ell}\mapsto o$ involves multiple nonlinear layers (attention, MLPs, layer norms). We denote this as:
$o=F_{\ell\to L}(h_{\ell}+\alpha v)$ (5)
where $F_{\ell\to L}$ is the composition of layers $\ell+1$ through $L$.
3.4 Assumptions
We make the following assumptions explicit.
A1 (Smoothness) The functions $g_{\ell}$ and $F_{\ell\to L}$ are differentiable almost everywhere, enabling local linear approximation via Jacobians $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}(x)$.
A2 (Identifiable prompts) The prompt distribution $p(x)$ has sufficient variability to probe different aspects of the latent persona $z$. Formally, the support of $p(x)$ is rich enough that different persona values induce distinguishable activation patterns $h_{\ell}(x)$.
A3 (Non-degeneracy) The Jacobian $J_{\ell}(x)$ has rank at least $k\geq 1$ for typical $x\sim p(x)$, meaning steering can affect outputs. This excludes pathological cases where all perturbations to $h_{\ell}$ are ignored by subsequent layers.
We do not assume: (i) statistical independence between $z$ and $x$ (confounding is allowed), (ii) linearity of $g_{\ell}$ in $z$ or (iii) uniqueness of the latent representation without additional structure. These assumptions are mild and generally satisfied by standard transformer architectures under typical operating conditions.
4 Definitions and Identifiability
4.1 Identifiability
Definition 1 (Parameter Identifiability)
A parameter $\theta$ in a statistical model $p(y\mid x;\theta)$ is identifiable if for any $\theta^{\prime}\neq\theta$, there exists a distribution over observations $(x,y)$ such that $p(y\mid x;\theta)\neq p(y\mid x;\theta^{\prime})$.
In our setting, the parameter is the steering vector $v\in\mathbb{R}^{d}$. We say $v$ is identifiable if no other vector $v^{\prime}\neq v$ (up to scaling) produces the same distribution over observable outputs across all prompts and steering strengths.
Definition 2 (Observational Equivalence)
Two steering vectors $v$ and $v^{\prime}$ are observationally equivalent in regime $\mathcal{R}$ if they produce identical distributions over all quantities observable in $\mathcal{R}$. For Regime 2 (white-box single-layer access):
$v\sim_{\mathcal{R}}v^{\prime}$ $\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)$ (6)
$\quad=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
4.2 Symmetries and Gauge Freedom
Scaling ambiguity
For any $c\neq 0$, the vectors $v$ and $cv$ produce outputs that differ only by a rescaling of $\alpha$. This is unavoidable; we consider $v$ and $cv$ as the same direction.
Null space ambiguity
If $v_{0}\in\ker(J_{\ell})$ (i.e., $J_{\ell}v_{0}=0$), then adding $v_{0}$ to any steering vector does not change the linearized output. Under linear approximation, $v$ and $v+v_{0}$ are observationally equivalent.
5 Main Results
We now state our main theoretical result, proving that persona vectors are fundamentally non-identifiable under standard observational regimes.
Proposition 1
Under Assumptions A1–A3 within the local linear approximation, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable. Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent under linearization.
Proof Sketch
We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in (Appendix B.2).
Null-space ambiguity (primary mechanism). Under the local linear approximation (Equation 4), the steered output is
$o(x,v,\alpha)\approx o_{0}(x)+\alpha J_{\ell}(x)v.$ (7)
For any $v_{0}\in\ker(J_{\ell})$, i.e., $J_{\ell}v_{0}=0$, define $v^{\prime}=v+v_{0}$. Then
$o(x,v^{\prime},\alpha)$ $\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})$ (8)
$=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha),$
Since $\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$, any case where $\mathrm{rank}(J_{\ell})<d$ implies a non-trivial null space. In overparameterized language models the effective rank of $J_{\ell}$ is often smaller than $d$ (Oymak et al., 2019), yielding $\dim(\ker(J_{\ell}))\geq 1$ in practice. Hence infinitely many such $v_{0}$ exist, producing observationally equivalent vectors $v^{\prime}\not\propto v$.
Remark (Scope of linearization)
Proposition 1 is proved under the local linear approximation (Equation 4), which holds exactly for piecewise-linear activations (ReLU MLPs on activation-region interiors) and as a first-order approximation for smooth nonlinearities (GeLU, attention softmax). For small steering strengths $\alpha$, the nonlinear residual is $O(\alpha^{2})$ and empirically negligible. Extending the result to the exact nonlinear regime $o=F_{\ell\to L}(h_{\ell}+\alpha v)$ would require characterizing the Hessian structure of $F_{\ell\to L}$ and is left for future work.
Corollary 1.1 (Null-space equivalence)
Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$ since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space. A complementary logit-level analysis in Appendix A further examines output distribution preservation beyond the semantic probes used in the main experiments.
6.1 Experimental Setup
Models and Layers
We evaluate two open-weight instruction-tuned language models to test generality across architectures and model scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$) (Yang et al., 2024; Meta AI, 2024). For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
Persona traits
We test three semantic traits spanning distinct dimensions: Formality, Politeness and Humor. This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
Steering vector extraction
For each trait, we construct 50 contrastive prompt pairs designed to elicit contrasting persona values. For example, formality pairs contrast "Write a professional and formal message about [topic]" versus "Write a casual and informal message about [topic]." For each prompt pair, we extract the hidden representation at layer $\ell$ for the final token position (Chen et al., 2025). The baseline steering vector is computed as the mean difference:
$v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})]$ (9)
Semantic probes
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous. Direct distributional evidence is provided by the logit-level analysis in Appendix A.
Orthogonality test methodology
To test Proposition 1’s prediction that $v$ and $v+v_{\perp}$ (where $v_{\perp}$ is orthogonal) produce observationally equivalent outputs, we implement the following procedure:
  1. 1.
    Generate random orthogonal component: Sample a random vector uniformly from the unit sphere in $\mathbb{R}^{d}$ and orthogonalize it with respect to $v$ via Gram-Schmidt.
  2. 2.
    Construct perturbed vector: Form $v^{\prime}=v+\alpha v_{\perp}$ where $\alpha$ is chosen such that $|v^{\prime}|\approx|v|$.
  3. 3.
    Generate steered outputs: For each of 100 held-out test prompts, generate text with steering vectors $v$ and $v^{\prime}$ at strength $\alpha=1.0$, producing 10 samples per prompt per vector.
  4. 4.
    Compute semantic equivalence: Measure Cohen’s $d$ effect size and Pearson correlation between semantic scores $\phi(o_{v})$ and $\phi(o_{v^{\prime}})$.
  5. 5.
    Repeat across seeds: Repeat for multiple random orthogonal seeds to assess robustness.
If $v$ and $v^{\prime}$ are observationally equivalent as predicted by Proposition 1, we expect Cohen’s $d<0.2$ (negligible effect) and high correlation between semantic scores.
Sample size design
We conduct the orthogonality test with both $n=5$ and $n=10$ random orthogonal seeds per trait for Qwen2.5-3B and Llama-3.1-8B. This allows us to assess statistical stability across sample sizes and models.
Scale invariance test
To verify that observational equivalence holds across different steering strengths, we additionally evaluate the formality trait at four $\alpha$ values: 0.0,0.5,1.0,2.0{0.0,0.5,1.0,2.0} for both models. For each $\alpha$, we measure semantic scores under steering with $v$ versus $v+v_{\perp}$ and plot response curves. If equivalence is scale-invariant, the curves should track closely across all $\alpha$ with overlapping confidence bands.
6.2 Orthogonal Perturbation Test Results
Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds). Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$. We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
Table 1: Empirical validation of non-identifiability across models, traits, and sample sizes. Cohen’s $d$ measures the effect size between $v$ and $v+v_{\perp}$ (smaller = more equivalent). All results show negligible differences ($d<0.2$), supporting observational equivalence.
Model Trait Seeds Cohen’s $d$ Correlation Perp-Only
Qwen2.5-3B-Instruct Formality $n=5$ $0.144\pm 0.095$ $0.210\pm 0.113$ $98.8\%$
$n=10$ $0.075\pm 0.058$ $0.285\pm 0.087$ $100.4\%$
Politeness $n=5$ $0.112\pm 0.077$ $0.319\pm 0.070$ $101.5\%$
$n=10$ $0.092\pm 0.069$ $0.414\pm 0.083$ $100.6\%$
Humor $n=5$ $0.103\pm 0.077$ $0.276\pm 0.177$ $99.7\%$
$n=10$ $072\pm 0.061 0.0$$044\pm 0.097$ $100.5\%$
Llama-3.1-8B-Instruct Formality $n=5$ $0.052\pm 0.029$ $0.164\pm 0.044$ $95.6\%$
$n=10$ $0.096\pm 0.068$ $0.192\pm 0.085$ $96.8\%$
Politeness $n=5$ $0.074\pm 0.041$ $0.324\pm 0.058$ $101.2\%$
$n=10$ $0.085\pm 0.043$ $0.347\pm 0.077$ $100.4\%$
Humor $n=5$ $159\pm 0.109 0.0$$032\pm 0.116$ $98.4\%$
$n=10$ $119\pm 0.119 0.0$$016\pm 0.104$ $95.9\%$
With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range. Near-zero correlations (e.g., $0.044$, 0.032-0.032) reflect probe ceiling effects rather than independence; Cohen’s $d$ is the more reliable metric in these cases. The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact. Values marginally above $100\%$ reflect probe ceiling effects at the upper boundary of the scoring range.
Statistical stability across sample sizes The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models. For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$). For Llama-3.1-8B, we observe similar stability: formality ($0.052\to 0.096$), politeness ($0.074\to 0.085$), humor ($0.159\to 0.119$). This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
Cross-model consistency The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific. Despite differences in architecture, scale (3B vs. 8B parameters) and hidden dimension ($d=2048$ vs. $d=4096$), both models exhibit nearly identical patterns of non-identifiability. The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
Visualizing orthogonal component equivalence Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds. The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
(a) Qwen2.5-3B-Instruct
(b) Llama-3.1-8B-Instruct
Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 indicate equivalence.
[…]
6.3 Scale Invariance Analysis
To verify that observational equivalence holds across different steering strengths, we evaluate the formality trait at four steering magnitudes $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for both models. Figure 3 shows the response curves for the extracted vector $v$ and the observationally equivalent vector $v+v_{\perp}$.
(a) Qwen2.5-3B-Instruct
(b) Llama-3.1-8B-Instruct
Figure 3: Scale invariance of observational equivalence. Formality scores across steering strengths $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for the extracted vector $v$ (blue circles) and the perturbed vector $v+v_{\perp}$ (orange squares).
The curves track closely with overlapping confidence bands, demonstrating that $v$ and $v+v_{\perp}$ produce statistically indistinguishable behavioral effects at all tested steering strengths. For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation). This proves that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
6.4 Robustness of Non-Identifiability Under Distribution Shift
We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability. We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing) and safety-style (refusal patterns, ethical dilemmas). For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
Table 2 presents Cohen’s $d$ effect sizes across all trait-environment combinations for both models. Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$). Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across prompt contexts (in-dis: in-distribution, top: topic shift, gen: genre shift, saf: safety-style).
Model Trait In-dis Top Gen Saf
Qwen-3B Formality 0.44 0.25 0.29 0.21
Politeness 0.22 0.24 0.52 0.41
Humor 0.42 0.51 0.24 0.51
Llama-8B Formality 0.30 0.54 0.30 0.21
Politeness 0.37 0.27 0.51 0.63
Humor 0.29 0.56 0.30 0.40
Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models). Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range. Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama). Figure 4 visualizes these patterns through environment-dependent effect sizes, showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
Figure 4: Environment-dependent effect sizes across prompt contexts.
Implications
The persistence of small effect sizes across diverse prompt distributions indicates that non-identifiability is not an artifact of in-distribution prompts but a structural geometric property. Cases where $d$ approaches 0.50.50.60.6 (e.g., politeness under genre or safety-style shifts) suggest that certain trait–context combinations impose partial constraints without eliminating null-space ambiguity since $\ker(J)$ is determined by model weights rather than the prompt distribution. Although effect sizes increase under distribution shift (reaching $d=0.63$ in some settings), distributional diversity alone does not eliminate observational equivalence.
7 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human evaluation, providing a conservative test of observational equivalence.
Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization or invariance objectives (IRM) can recover identifiable representations. Empirical validation of these approaches remains an important direction for future work. Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
8 Future Work
Our results establish the necessity of validation beyond input-output behavior. Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components and characterizing how null-space dimensionality scales with model size.
Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability. Human evaluation would provide stronger validation than our lexical heuristics and direct measurement of null-space dimensionality would quantify our theoretical predictions.
9 Conclusion
We have established that steering vectors are fundamentally non-identifiable under the local linear approximation, with infinitely many distinct directions producing equivalent behavioral effects. This non-identifiability persists across models, traits and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may fail to yield interpretable representations. Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone. More broadly, when internal interventions admit large equivalence classes under the model’s local geometry, consistent output shifts do not imply that a discovered direction corresponds to a canonical or causally privileged feature. Stronger interpretability claims therefore require structural constraints or additional assumptions that break these symmetries and distinguish meaningful factors from geometrically arbitrary ones.
Acknowledgments
We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments. We also acknowledge the HuggingFace platform and maintainers of the open-source models used in this work. Generative AI tools were used for language editing and code assistance; all technical content was reviewed and verified by the authors.
Appendix A Logit-Level Equivalence Test
To complement our semantic evaluations, we perform a direct logit-level equivalence test. For each model and trait, we compare the next-token logit vectors produced by steering with $v$ and with $v+v_{\perp}$, where $v_{\perp}$ is a norm-matched direction orthogonal to $v$. We measure the $\ell_{2}$ distance between logit vectors and normalize by the corresponding distance induced by a norm-matched random direction:
$R=\frac{\|\ell(v)-\ell(v+v_{\perp})\|_{2}}{\|\ell(v)-\ell(v_{\text{rand}})\|_{2}}$ (10)
Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions. In all tested models and traits, this condition holds. Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions. The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$ and top-10 token overlap ranges from $86\%$ to $96\%$.
These results provide direct evidence that equivalence is not an artifact of semantic measurement. While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes. Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement and top-$k$ overlap for all models and traits.
Table 3: Logit-level equivalence statistics. $R$ denotes the ratio $v_{\perp}$ / random. Agreement% refers to token agreement, and Overlap% indicates top-10 overlap.
Model Trait $R$ Agreement% Overlap%
Qwen-3B Humor 0.90 81% 86%
Formality 0.76 94% 96%
Politeness 0.75 93% 96%
Llama-8B Humor 0.64 96% 90%
Formality 0.49 93% 95%
Politeness 0.65 91% 94%
Appendix B Detailed Proof of Proposition 1
B.1 Null-Space Ambiguity (Constructive Proof)
Setup Consider the local linear approximation of the steering effect:
$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (11)
where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian.
Step 1: Null space characterization
Define the null space of $J_{\ell}$ as:
$\mathcal{N}=\ker(J_{\ell})=\{v_{0}\in\mathbb{R}^{d}:J_{\ell}v_{0}=0\}$ (12)
By the rank-nullity theorem:
$\dim(\mathcal{N})=d-\operatorname{rank}(J_{\ell})$ (13)
Step 2: Rank Bound
The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ has maximum possible rank:
$\text{rank}(J_{\ell})\leq\min(V,d)$ (14)
In modern language models, the hidden dimension is typically $d\approx 4000$ while the vocabulary size is $V\approx 50000$. Therefore, the Jacobian satisfies $\max\mathrm{rank}(J_{\ell})=d$.
Step 3: Effective rank is much lower
Output distributions lie on a low-dimensional manifold. The effective rank satisfies:
$\text{rank}_{\epsilon}(J_{\ell})=\#\{\sigma_{i}:\sigma_{i}>\epsilon\cdot\sigma_{\max}\}\ll d$ (15)
where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\epsilon$ is a threshold (e.g., $10^{-4}$).
Intuition: Empirical evidence suggests that in overparameterized LLMs the effective rank of $J_{\ell}$ may be substantially lower than $d$, consistent with low dimensional structure observed in related weight space studies (Li et al., 2018; Maennel et al., 2018) and concentration of output sensitivity in a limited number of directions (Ghorbani et al., 2019). Consequently, $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is often positive in practice. However, this empirical observation is not required for the proof; the argument establishes non identifiability whenever $\dim(\mathcal{N})\geq 1$.
Step 4: Constructing equivalent vectors
For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define
$v^{\prime}=v+v_{0}.$ (16)
Then for all $x$ and all $\alpha$:
$J_{\ell}(x)v^{\prime}$ $=J_{\ell}(x)(v+v_{0})$ (17)
$=J_{\ell}(x)v+J_{\ell}(x)v_{0}$
$=J_{\ell}(x)v.$
Therefore:
$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (18)
$=o(x,0,0)+\alpha J_{\ell}(x)v$
$\approx o(x,v,\alpha).$
Step 5: Infinitely many distinct solutions
Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$:
$v^{\prime}_{\beta}=v+\beta v_{0}$ (19)
generates infinitely many observationally equivalent vectors. Furthermore, if $\beta$ is chosen such that $v^{\prime}_{\beta}\not\propto v$ (which is always possible unless $v\propto v_{0}$), these vectors are geometrically distinct.
Step 6: Non-proportionality
To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$. This fails only if:
$\beta v_{0}=(c-1)v$ (20)
which requires $v\in\text{span}(v_{0})$. Since $v_{0}$ is an arbitrary element of a $\dim(\mathcal{N})$-dimensional space and $v$ is fixed, this occurs with probability zero. Therefore, for generic $v$ and generic $v_{0}\in\mathcal{N}$, we have $v^{\prime}_{\beta}\not\propto v$ for almost all $\beta$.
Conclusion (Null-space mechanism)
Under the local linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
B.2 Weight Space Gauge Symmetry (Supplementary Observation)
While Proposition 1 fixes model weights and asks whether v is recoverable from behavior, this section offers a complementary perspective: even relaxing the weight-fixing assumption, non-identifiability persists via linear reparameterization symmetries.
Setup: We record a supplementary observation about gauge symmetry in neural network parameterization. This is distinct from the null-space non-identifiability of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs can produce identical input–output behavior. The answer is yes, via the linear reparameterization below. This symmetry is not realizable as a steering operation on a frozen model as it requires modifying $W_{\ell+1}$.
Step 1: Representation reparameterization
Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Define the reparameterized representation:
$h^{\prime}_{\ell}(x)=T\bigl(h_{\ell}(x)\bigr).$ (21)
If the subsequent layers can be rewritten as:
$F_{\ell\to L}(h_{\ell})=F^{\prime}_{\ell\to L}(T(h_{\ell}))$ (22)
then $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable.
Step 2: Linear reparameterizations
Consider linear transformations $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible. For layer $\ell+1$ with weight matrix $W_{\ell+1}\in\mathbb{R}^{d\times d}$ and bias $b_{\ell+1}$:
Original computation:
$h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ (23)
Reparameterized computation:
$h^{\prime}_{\ell+1}=\sigma(W^{\prime}_{\ell+1}h^{\prime}_{\ell}+b^{\prime}_{\ell+1})$ (24)
where $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$ and $b^{\prime}_{\ell+1}=b_{\ell+1}$.
Note: The bias remains invariant under linear reparameterization through the origin. For more general affine transformations $T(h)=Ah+c$ with translation $c\neq 0$, the bias would also transform as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$. We restrict to origin-preserving transformations for simplicity, as these suffice to establish non-identifiability. Then:
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (25)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}.$
Step 3: Steering under reparameterization
Original steering:
$\tilde{h}_{\ell}=h_{\ell}+\alpha v$ (26)
$\tilde{h}_{\ell+1}=\sigma(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1})$ (27)
Reparameterized steering with $v^{\prime}=Av$:
$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (28)
$=Ah_{\ell}+\alpha Av$
$=A(h_{\ell}+\alpha v)$
$=A\tilde{h}_{\ell}$.
$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$ (29)
$=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
$=\sigma\!\left(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
$=\tilde{h}_{\ell+1}.$
Step 4: Infinitely many reparameterizations
For any invertible matrix $A$ with $A\neq cI$ (i.e., not a scalar multiple of identity), we obtain $v^{\prime}=Av\not\propto v$. The space of such matrices has dimension $d^{2}-1$ (excluding scalar multiples), providing infinitely many distinct reparameterizations.
Conclusion (Weight-space gauge symmetry)
The symmetries of neural network parameterization imply the existence of infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs that yield identical input–output behavior.
Appendix C Intuitive Explanations
C.1 Geometric Intuition for Null-Space Ambiguity
Visual analogy:
Consider a simplified example where a 3D steering vector $v\in\mathbb{R}^{3}$ affects 2D outputs $o\in\mathbb{R}^{2}$ through a projection matrix $J\in\mathbb{R}^{2\times 3}$. By the rank-nullity theorem, the null space $\ker(J)$ is a 1D subspace since $\dim(\ker(J))=3-\text{rank}(J)=3-2=1$.
[…]
Figure 5: Geometric intuition for null-space ambiguity. The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible; hence $J(v+\alpha v_{3})=Jv$.
For concreteness, suppose $J=[I_{2}\mid 0]$ projects onto the first two coordinates. Then $\ker(J)=\text{span}\{(0,0,1)\}$ is the $v_{3}$ axis. The key insight generalizes: for any projection matrix $J$, directions in $\ker(J)$ are invisible to outputs. Figure 5 illustrates this geometric intuition.
Any vector of the form $v^{\prime}=v+\alpha v_{3}$ for $\alpha\in\mathbb{R}$ produces identical outputs:
$J(v+\alpha v_{3})$ $=Jv+\alpha Jv_{3}$ (30)
$=Jv+\alpha\cdot 0\quad(v_{3}\in\ker(J))$
$=Jv.$
Since $\alpha$ can take infinitely many values and $v^{\prime}\not\propto v$ for generic choices of $\alpha$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
C.2 Why More Data Doesn’t Help
Common misconception: "If we collect more steering examples, we can pin down the unique vector."
Why this fails:
Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$. More fundamentally, any $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v_{0}=0$ by definition since this is a property of the model weights and not the prompt distribution. Stacking observations across prompts therefore cannot remove directions from $\ker(J_{\ell})$. Each observation provides:
$o_{i}=J_{\ell}(x_{i})v+\eta_{i}$ (31)
Stacking these:
$\begin{bmatrix}o_{1}\\ o_{2}\\ \vdots\\ o_{N}\end{bmatrix}=\begin{bmatrix}J_{\ell}(x_{1})\\ J_{\ell}(x_{2})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}v+\begin{bmatrix}\eta_{1}\\ \eta_{2}\\ \vdots\\ \eta_{N}\end{bmatrix}$ (32)
The stacked Jacobian $J_{\text{stack}}\in\mathbb{R}^{(N\cdot V)\times d}$ has null space:
$\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (33)
[…]
Figure 6: Illustration of why prompt diversity does not resolve null-space ambiguity. Each prompt induces a Jacobian that measures the same effective subspace, leaving the null-space intersection unchanged.
Critical observation: If all $J_{\ell}(x_{i})$ share a common null space (e.g., all prompts probe similar aspects of the model), then:
$\ker(J_{\text{stack}})=\ker(J_{\ell}(x_{1}))\neq\{0\}$ (34)
This means adding more prompts does not reduce the null space when all prompts probe the same effective subspace—the ambiguity persists. Figure 6 illustrates this phenomenon.
C.3 Why Orthogonal Perturbations Preserve Semantics
Setup: We test steering with $v$ versus $v+v_{\perp}$, where $v_{\perp}\perp v$ and $\|v_{\perp}\|=1$. Empirically, adding a random orthogonal direction produces nearly equivalent semantic effects. Why?
Decomposing the perturbation
Any perturbation $v_{\perp}$ can be decomposed into observable and invisible components:
$v_{\perp}=v_{\perp,\text{row}}+v_{\perp,\text{null}}$ (35)
where $v_{\perp,\text{row}}\in\operatorname{row}(J)$ (observable through outputs) and $v_{\perp,\text{null}}\in\ker(J)$ (invisible to outputs).
For a random $v_{\perp}\perp v$, the expected fraction in the null space is:
$\mathbb{E}[\|v_{\perp,\text{null}}\|^{2}]\approx\frac{\dim(\ker(J))}{d}\approx 0.20{-}0.25$ (36)
This means $\sim$20–25% of the perturbation is automatically invisible to model outputs.
Observable effect is diffuse and unstructured
The actual output change from adding $v_{\perp}$ is:
$J(v+v_{\perp})=Jv+Jv_{\perp,\text{row}}$ (37)
While $Jv_{\perp,\text{row}}$ is not necessarily small in norm, it represents a random direction in the $\sim$3200-dimensional row space of $J$. In contrast, $Jv$ is a structured, semantically coherent steering effect (e.g., consistently shifting outputs toward “honesty” or “sycophancy”).
The key distinction is not magnitude but semantic coherence: $Jv$ produces aligned, directional changes across the vocabulary, while $Jv_{\perp,\text{row}}$ produces diffuse, incoherent perturbations that average out across tokens and do not systematically shift semantic meaning in any consistent direction.
Analogy: If $Jv$ is a strong wind blowing consistently north, then $Jv_{\perp,\text{row}}$ is turbulence—it may have comparable energy but lacks directional coherence, so the overall trajectory remains northward.
Appendix D Prompt Construction Details
We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
Table 4: Example contrastive prompt pairs for each semantic trait.
Trait Positive ($x^{+}$) Negative ($x^{-}$)
Formality Draft a professional email summarizing quarterly performance. Text a friend about how the quarter went.
Politeness Respond diplomatically to criticism from a colleague. Reply bluntly to someone who criticized you.
Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
Figure 7: Steering extraction and evaluation pipeline.
Figure 7 illustrates the pipeline used to extract and evaluate steering vectors. Contrastive prompt pairs are first used to compute a steering vector specific to a semantic trait. This vector is then applied to a separate set of held-out evaluation prompts to generate steered outputs.
Appendix E Mathematical Derivations
E.1 Dimension of Observational Equivalence Class
Setup: For steering vector $v\in\mathbb{R}^{d}$ and Jacobian $J\in\mathbb{R}^{V\times d}$ with rank $r$, consider:
$v]=\{v^{\prime}\in\mathbb{R}^{d}:Jv^{\prime}=Jv\}$ (38)
Question: How many degrees of freedom exist in the equivalence class of observationally equivalent steering vectors?
Analysis: The equivalence class is an affine subspace:
$v]=v+\ker(J)$ (39)
where $\ker(J)$ is a (dr)(d-r)-dimensional linear subspace.
Parameterization: Let $\{u_{1},\ldots,u_{d-r}\}$ be an orthonormal basis for $\ker(J)$. Then:
$v]=\left\{v+\sum_{i=1}^{d-r}\alpha_{i}u_{i}:\alpha_{i}\in\mathbb{R}\right\}$ (40)
The equivalence class has (dr)(d-r) degrees of freedom.
Scaling ambiguity: Additionally, $v$ and $cv$ produce equivalent outputs (up to rescaling $\alpha$). The equivalence class modulo scaling is a (dr)(d-r)-dimensional projective space.
Illustrative example. Suppose a model has hidden size $d=4096$ and an effective Jacobian rank $r\approx 3100$. Then the equivalence class dimension is $d-r\approx 996$.
Interpretation: For every identifiable parameter (row-space direction), there are $996/3100\approx 0.32$ unidentifiable parameters (null-space directions). The under-determination ratio is approximately $1:3$.
E.2 Fisher Information and Cramér-Rao Bound
Setup: Consider the statistical model:
$o(x;v,\alpha)\sim p(o|x,v,\alpha)$ (41)
where $v$ is the parameter to estimate.
Question: Can we achieve better identifiability by collecting more samples?
Fisher Information Matrix: For the linear Gaussian model $o=Jv+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}I)$:
$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (42)
$\qquad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
$=\frac{1}{\sigma^{2}}J^{\top}J.$
Cramér-Rao Lower Bound: The covariance of any unbiased estimator $\hat{v}$ satisfies:
$\text{Cov}(\hat{v})\succeq\mathcal{I}(v)^{-1}=\sigma^{2}(J^{\top}J)^{+}$ (43)
where $(J^{\top}J)^{+}$ is the Moore-Penrose pseudo-inverse.
Implications: For null-space directions $u_{i}\in\ker(J)$, the Fisher information is degenerate:
$u_{i}^{\top}(J^{\top}J)u_{i}=0$ (44)
Therefore, the variance of any unbiased estimator along null-space directions is unbounded:
$\text{Var}(u_{i}^{\top}\hat{v})\geq u_{i}^{\top}\mathcal{I}(v)^{-1}u_{i}=\infty$ (45)
Conclusion: The Cramér-Rao bound is infinite for null-space components, confirming that no finite amount of data can resolve the ambiguity. Non-identifiability is fundamental, not a small-sample problem. More data cannot help because the information geometry has infinite uncertainty in null-space directions.